OpenStudy (anonymous):

Will Give Medal. Is -7 +9 = -9 + 7 true, false, or open? Is 4x - 3 = 19 true, false, or open?

OpenStudy (anonymous):

-7+9=2 -9+7=2 so its false

OpenStudy (anonymous):

4x-3=19 4x=22 x=5.5

OpenStudy (carolinar7):

open I think for that one

OpenStudy (anonymous):

i think so too